describe and correct the error in solving the linear system. 9x + 8y = 11 and 7x + 6y = 9
Answer by stanbon(75887)

9x + 8y = 11
7x + 6y = 9
-----
Modify for elimination:
63x + 56y = 77
63x + 53y = 81
==========
Subtract and solve for "y":
3y = -4
y = -4/3
------------------
Solve for "x":
7x + 6y = 9
7x + 6(-4/3) = (27/3)
-----
7x = 51/3
x = 51/21
--------------------
Cheers,
Stan H.
====================